Given a 9, b= 7, and C=15, use the law of cosines to solve the triangle for the valve b А c B
Answer #1

Given

a = 9

b = 7

c = 15

Apply law of cosine

c2 = a2 + b2 - 2*a*b*cos( C )

=> c2 - a2 - b2 = - 2*a*b*cos( C )

=> [ c2 - a2 - b2 ] / ( - 2*a*b ) = cos( C )

Put the values of a , b and c

=> [ 152 - 92 - 72 ] / ( - 2*9*7 ) = cos( C )

=> [ 225 - 81 - 49 ] / ( - 126 ) = cos( C )

=> - 95 / 126 = cos( C )

=> C = arccos( - 95 / 126 )

=> C = 138.93 degree